# Sensor-fusion based Prognostics Framework for Complex Engineering Systems Exhibiting Multiple Failure Modes

### v2

@@ 5 Online diagnosis and Prognostics
This section details the steps of the online part of our multi-failure mode prognostics framework. In this part, a fielded unit is degrading in real-time, and our goal is to predict the time remaining before the unit fails. We begin by showing how MFPCA is used to fuse the signals from multiple sensors into a parsimonious set of features. Then, we demonstrate how we utilize MFPCA to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
1
Input : Smoothed signals from training & test units, Failure labels of training data from Algorithm 1, Set of informative sensors $\mathcal{P}_{k}$ from Algorithm 1
Output : Active failure mode in the test unit $k^{*}$, Remaining Useful Life (RUL) of the test unit $RUL_{test}$
2
/* Diagnostic Model Training */
3 $\zeta_{(i),h}\leftarrow$ MFPCA on training units (Eq. 10);
4 Train a KNN classifier using $\zeta_{(i),h}$;
5
/* Diagnostic Model Testing */
6 $\zeta_{test,h}\leftarrow$ MFPCA on test unit;
7 $k^{*}\leftarrow$ Mode of labels in KNN of $\zeta_{test,h}$;
8
/* Prognostic Model Training */
9 $\zeta^{k^{*}}_{(i),h}\leftarrow$ MFPCA: selected sensors (Eq. 14);
10 $c_{0}^{k^{*}},\dots,c_{h}^{k^{*}}\leftarrow$ Obtain $\operatorname*{arg\,min}$ in Eq. 17;
11
/* Prognostic Model Testing */
12 $\zeta^{k^{*}}_{test,h}\leftarrow$ MFPCA on test unit (Eq. 14);
13 $RUL_{test}\leftarrow$ Equation 18;
Algorithm 2 Online Diagnosis and Prognostics
@@ 5.1 Sensor Fusion: MFPCA
Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse. Now assume that we have monitored an online unit (operating in the field) for a duration given by the time interval $[0,t^{*}]$. Since we are interested in predicting the remaining lifetime, we only use signals from units with failure times greater than $t^{*}$. We define this set as the “$t^{*}-$training” set, whose indices we denote by the set $\mathcal{I}_{t^{*}}$, which has cardinality $N_{t^{*}}$. Furthermore, we smooth the degradation signals using the ’rloess’ method as described in [40]. This method uses weighted least squares with a 2nd-order polynomial for smoothing. The weights come from a bisquare weight function requiring a specified bandwidth parameter. This method also identifies outliers and assigns them weights of zero to improve signal mean estimation. To perform MFPCA, we model the $i$th smoothed multivariate signal evaluated at time $t$ as $\boldsymbol{s}_{i}(t)=\boldsymbol{u}_{i}(t)+\boldsymbol{\epsilon}_{i}(t)$, where $\boldsymbol{u}_{i}(t)\in\mathcal{R}^{\mathcal{P}}$ is a random multivariate function with mean $\boldsymbol{\mu}(t)$ and covariance function $\mathcal{C}(t,t^{\prime})$, $\boldsymbol{\epsilon}_{i}(t)\in\mathcal{R}^{\mathcal{P}}$, and $t\in[0,t^{*}]$ are i.i.d. zero-mean error vectors with covariance $\Sigma_{\boldsymbol{\epsilon}}$. We assume $\boldsymbol{u}_{i}(t)$ and $\boldsymbol{\epsilon}_{i}(t)$ are independent of each other. Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by equation 10 where the MFPC-scores $\zeta$’s are computed according to equation 11.
@@ 5.2 Failure Mode Diagnosis: K-nearest Neighbors
In order to diagnose the failure mode of a degrading unit at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset. Next, $K$ nearest neighbors (KNN) is applied to the extracted MFPC scores to identify the active failure mode. Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC scores. Given this diagnosis, we now develop a model for predicting the unit RUL.
@@ 5.3.1 Time-Varying Functional Regression
When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (FPC scores). The predictor trajectories up to the current time are represented by time-varying FPC scores, which are continuously updated as time progresses and are considered to be time-varying predictor variables for the functional regression model. This is referred to as time-varying functional regression. It was first proposed by [41] and has been used for prognostics in [42, 13].
In our problem setting, let $k^{*}$ denote the diagnosed failure mode and $\mathcal{I}_{t^{*}}^{k^{*}}\subset\mathcal{I}_{t^{*}}$ denote the set of indices for training units that failed due to failure mode $k^{*}$, but survived up to time $t^{*}$. Furthermore, let $\mathcal{P}_{k^{*}}\subset\{1,2,...,\mathcal{P}\}$ denote the set of sensors selected (identified as most informative) for failure mode $k^{*}$. To predict RUL, we fit a functional regression model using least squares as shown in equation (12), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
@@
We expand the coefficient function using the eigenfunctions obtained from applying MFPCA on the training set $\mathcal{I}_{t^{*}}^{k^{*}}$. The regression coefficient function can now be expanded as shown in equation (13), where $H_{k^{*}}$ denotes the number of MFPC-scores that were retained for failure mode $k^{*}$.
@@
Likewise, we expand the multivariate signals using the eigenfunctions as shown in equation (14).
@@
Since the eigenfunctions are orthogonal, substituting equations (13) and (14) into equation (12) yields the following optimization problem,
@@
Weighted Version: We note that the offline sensor selection algorithm may generate some outliers, which can impact the accuracy of the regression problem. These outliers become more apparent after identifying the active failure mode and the MFPC scores to estimate the time-varying functional regression model. To address this issue, we use weighted least squares to estimate the time-varying functional regression mode. To derive the weights, we first measure the Euclidean distance between the MFPC score corresponding to a unit $i$ and the centroid of the cluster of the active failure mode. We then use the reciprocal of that distance to define the corresponding weight for the unit, $w_{i}$, of that unit. If the unit’s failure is misclassified, $w_{i}$ would be large, and thus, unit $i$ is assigned a smaller weight in the regression problem. Next, We regress the $\ln{TTF}$ on the MFPC scores using a weighted linear regression shown in equation (17). The solution to this problem can be found using weighted least squares regression [43]. For variable selection, we include the lasso penalty in the optimization problem.
@@ 5.3.2 RUL prediction
As we have shown, MFPCA enables the transformation of the functional regression problem to a multiple linear regression problem, which enables us to solve the problem only by estimating $H_{k^{*}}+1$ parameters. After fitting the regression coefficients, we extract the MFPC scores for the test signal $\zeta_{test,1}^{k^{*}},\ldots,\zeta_{test,H_{k^{*}}}^{k^{*}}$. These scores are then used to predict the RUL at time $t^{*}$ using equation (18).
${RUL}_{test}=\exp\left(c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\right)-t^{*}$ (18)
@@ 6 Case Study 1: Simulated Dataset
We present a case study using a simulated data set to evaluate the performance of our methodology in terms of its ability to cluster the training data, effectiveness in selecting informative sensors for each failure mode, and accuracy in predicting the RUL. Furthermore, we test the robustness of our methodology to varying degrees of signal noise. In section VI-A, we describe the simulation model. In section VI-B, we display the simulated data and evaluate the performance of the MGR-ASGL model at clustering and sensor selection. Finally, we evaluate the prediction accuracy of our online prognostics approach in section VI-C.
@@ 6.1 Simulation Model
We utilize a modified version of the simulation model used in [13] to generate degradation signals for a set of engineering systems. We assume that each system is monitored by 20 sensors and exhibits one of two possible failure modes that we refer to as failure mode 1 (FM 1) and failure mode 2 (FM 2). We generate data for 200 systems for each failure mode. Out of the 200 systems, we use 160 systems for model training and 40 systems for testing model performance. The simulated data consists of the sensor signals for each system and the corresponding failure time, which is utilized later as the ground truth to evaluate prediction accuracy.
The signals of the underlying degradation process for system $i$ exhibiting failure mode $k$ are generated according to equation (19). The signal is parameterized by the random coefficient $\theta_{i}^{k}\sim N\left(\mu_{k},0.1^{2}\right)$, which is used to control the degradation rate. We let $\mu_{1}=1$, and $\mu_{2}=0.8$ be the degradation rates for FM 1 and FM 2, respectively. The TTF is computed as the first time the degradation signal $s_{i}^{k}(t)$ crosses a soft failure threshold $D_{k}$. We set the failure thresholds for FM 1 and FM 2 as $D_{1}=2$ and $D_{2}=1.5$, respectively.
@@
This paragraph describes the procedure for simulating degradation signals for each sensor. Here we consider a practical scenario where only a subset of sensors is correlated with the underlying degradation process associated with a given failure mode. These sensors are the informative sensors and their signal trends reflect the severity and progression of the degradation process. Note that sensors that do not exhibit any consistent trends leading up to a failure event are deemed noninformative. We define $\mathcal{J}_{k}$ as the index set of informative sensors for failure mode $k$, and $\mathcal{N}_{k}$ as the set of noninformative sensors such that $\mathcal{N}_{k}\cap\mathcal{J}_{k}=\{\}$ (null set). Out of the 20 sensors, we set $\mathcal{J}_{1}=\{5,12,16,19\}$ as the set of informative sensors for FM 1 and $\mathcal{J}_{2}=\{3,7,9,19\}$ as the set of informative sensors for FM 2. The $p^{th}$ sensor signal for the $i^{th}$ system exhibiting failure mode $k$ is generated using equation (20).
@@
In this equation, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the signal-to-noise for sensor $p$ and failure mode $k$. To account for the correlation between the signal trends of the informative sensors and the underlying degradation process, we generate $\theta_{i,p}^{k}$ from the following conditional distribution $\theta_{i,p}^{k}\mid\theta_{i}^{k}\sim N\left(\mu_{k}\left(1-\sqrt{1-\rho_{p}^{k}}\right),(0.1)^{2}\right)$ such that the correlation, $\rho_{p}^{k}$, is uniformly sampled from the interval $[0.80,0.99]$. The process for generating signals for the noninformative sensors is similar to that of the informative sensors, except the interval from which we sample $\rho_{p}^{k}$ is $[0.1,0.6]$. Finally, we randomly assign a sign to $\theta_{i,p}^{k}$ to allow for signals that increase or decrease in response to degradation.
@@ 6.2 Failure Mode Labeling and Sensor Selection
We analyze the ability of the MGR-ASGL model to accurately cluster the degradation signals into one of the two possible failure modes along with selecting the subset of informative sensors. We start by utilizing CA-FPCA as mentioned in 4.1 to extract features for each sensor separately. To determine the number of CA-FPC scores to retain, we use the 95% fraction of variance explained (FVE) criterion. For each cluster, we select the first $q_{p}^{k}$ CA-FPC-scores whose eigenvalues sum to at least 0.95. Since the number of CA FPC scores to retain can vary for each cluster, we retain the maximum of $q_{p}^{k}$ over both failure modes. For the remainder of this paper, we use 95% FVE criterion for determining the number of FPC scores, CA-FPC scores, and MFPC scores to retain. The CA-FPC scores serve as inputs to the MGR-ASGL model. Since this model has two tuning parameters, we utilize 5-fold cross-validation with $\lambda\in\{0.0050,0.0258,0.0466,...,0.3792,0.4000\}$ and $\alpha\in\{0.00,0.25,0.50,0.75,1.00\}$. We use minimum mean-squared error (minMSE) criterion, which selects the tuning parameters that minimize the mean-squared prediction error averaged over all folds. Table 1 displays the tuning parameter selection using both criteria over all three datasets. For each dataset and criterion, we fit the MGR-ASGL model using these tuning parameters.
@@
We then evaluate the performance of the MGR-ASGL model on the three datasets. While clustering is an unsupervised learning task, we know the ground truth regarding the failure modes. Therefore, we evaluate the ability of the clustering algorithm to match the ground truth. Furthermore, we have predetermined which sensors are most informative and we would like to determine whether or not the MGR-ASGL selects these sensors. The results of this performance evaluation are shown in Table 2.

### v3

@@ 5 Online Diagnosis and Prognostics
In this section, a subsystem(s) in deep-space habitat is considered to be degrading in real-time. Our goal is to predict the time remaining before the unit fails. We begin by showing how to fuse the signals from multiple sensors into a parsimonious set of features. Then, we demonstrate how we utilize the fused sensor information to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
@@ 5.1 Sensor Fusion: MFPCA
Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse. Now assume that we have monitored the degrading subsystem(s) for a duration given by the time interval $[0,t^{*}]$. Since we are interested in predicting the remaining lifetime, we only use signals from units with failure times greater than $t^{*}$. We define this set as the “$t^{*}-$training” set, whose indices we denote by the set $\mathcal{I}_{t^{*}}$, which has cardinality $N_{t^{*}}$. Furthermore, we smooth the degradation signals using the ’rloess’ method as described in [52]. This method uses weighted least squares with a 2nd-order polynomial for smoothing. The weights come from a bisquare weight function requiring a specified bandwidth parameter. This method also identifies outliers and assigns them weights of zero to improve signal mean estimation. To perform MFPCA, we model the $i$th smoothed multivariate signal evaluated at time $t\in[0,t^{*}]$ as having mean $\boldsymbol{\mu}(t)\in\mathcal{R}^{\mathcal{P}}$ and $\mathcal{P}\times\mathcal{P}$ block covariance function $\mathcal{C}(t,t^{\prime})$, where $\mathcal{C}_{j,j^{\prime}}(t,t^{\prime})=cov(v_{j}(t),v_{j^{\prime}}(t^{\prime}))$. Using Mercer’s theorem [49], we can represent the covariance function as $\mathcal{C}(t,t^{\prime})=\sum_{h=1}^{\infty}\eta_{h}\boldsymbol{\varpi}_{h}(t)\boldsymbol{\varpi}_{h}(t^{\prime})^{T}$, where $(\eta_{h},\boldsymbol{\varpi}_{h}(t))$ are the $h$th eigenvalue-eigenfunction pair and $\eta_{1}\geq\eta_{2}\geq\dots$. Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by Eq 13 where the MFPC-scores $\zeta$’s are computed according to Eq 14.
@@
In practice, we perform MFPCA by collecting the $N_{t^{*}}$ observations, each consisting of the $\mathcal{P}$ sensor signals concatenated into a single vector. Then we perform traditional PCA on the sample of observations. Furthermore, we standardize the extracted MFPC-scores so they have zero mean and unit variance.
@@ 5.2 Failure Mode Diagnosis: K-nearest Neighbors
In order to diagnose the failure mode of a degrading unit at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset. Next, $K$ nearest neighbors (KNN) is applied to the standardized MFPC-scores to identify the active failure mode. Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC-scores. Given this diagnosis, we now develop a model for predicting the unit RUL.
@@ 5.3.1 Time-Varying Functional Regression
When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (MFPC-scores). The predictor trajectories up to the current time are represented by time-varying MFPC-scores, which are continuously updated as time progresses and are considered to be time-varying predictor variables for the functional regression model. This is referred to as time-varying functional regression. It was first proposed by [53] and has been used for prognostics in [54, 20]. In our problem setting, let $k^{*}$ denote the diagnosed failure mode and $\mathcal{I}_{t^{*}}^{k^{*}}\subset\mathcal{I}_{t^{*}}$ denote the set of indices for training units that failed due to failure mode $k^{*}$, but survived up to time $t^{*}$. Furthermore, let $\mathcal{P}_{k^{*}}\subset\{1,2,...,\mathcal{P}\}$ denote the set of sensors selected (identified as most informative) for failure mode $k^{*}$. To predict RUL, we fit a functional regression model as shown in Eq (15), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
@@
We expand the coefficient function using the eigenfunctions obtained from applying MFPCA on the training set $\mathcal{I}_{t^{*}}^{k^{*}}$. The regression coefficient function can now be expanded as shown in Eq (16), where $H_{k^{*}}$ denotes the number of MFPC-scores that were retained for failure mode $k^{*}$.
@@
Likewise, we expand the multivariate signals using the eigenfunctions as shown in Eq (17).
@@
Since the eigenfunctions are orthogonal, substituting Eqs (16) and (17) into Eq (15) yields the following optimization problem,
@@
Weighted Version: We note that the offline sensor selection algorithm may generate some outliers, which can impact the accuracy of the regression problem. These outliers become more apparent after identifying the active failure mode and the MFPC-scores to estimate the time-varying functional regression model. To address this issue, we use weighted least squares to estimate the time-varying functional regression model. To derive the weights, we first measure the Euclidean distance between the MFPC score corresponding to a unit $i$ and the centroid of the cluster of the active failure mode. We then use the reciprocal of that distance to define the corresponding weight for the unit, $w_{i}$, of that unit. If the unit’s failure is misclassified, $w_{i}$ would be large, and thus, unit $i$ is assigned a smaller weight in the regression problem. Next, We regress the standardized $\ln{TTF}$ on the MFPC-scores using a penalized weighted linear regression shown in Eq (20). For variable selection, we include the lasso penalty in the optimization problem. To ensure a regression coefficient represents the significance of their respective predictor to the prediction of the failure time, Eq (20) should be read such that $\zeta_{i,h}^{k^{*}}$ is standardized.
@@ 5.3.2 RUL prediction
As we have shown, MFPCA enables the transformation of the functional regression problem into a multiple linear regression problem, which allows us to solve the problem by estimating only $H_{k^{*}}+1$ parameters. After fitting the regression coefficients, we extract the MFPC-scores for the test signal $\zeta_{test,1}^{k^{*}},\ldots,\zeta_{test,H_{k^{*}}}^{k^{*}}$. These MFPC scores are standardized using the sample mean and sample standard deviation of the MFPC scores from the training set. These scores are then used to predict the RUL at time $t^{*}$ using Eq (21).
$\begin{split}{RUL}_{test}=\exp&\bigg{(}\Big{[}c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\Big{]}\times\\ &\Big{[}(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}\Big{]}\\ &+(1/N)\sum_{i=1}^{N}\ln{TTF_{i}}\bigg{)}-t^{*}\end{split}$ (21)
@@ 6 Case Study 1: Simulated Dataset
We present a simulation study that mimics the operational and diagnostic challenges of a deep-space habitat. In such systems, critical subsystems like life support, power generation, etc., are monitored by numerous sensors, and degradation can occur through multiple unknown failure modes. Due to the scarcity of in-mission failure data and the infeasibility of performing full-scale failure testing in space, simulation-based validation is pursued in this case study. This study enables us to systematically evaluate our methodology’s ability to: (1) identify latent failure modes in unlabeled multivariate sensor data, (2) select informative sensors specific to each mode, and (3) accurately predict remaining useful life (RUL).
To emulate the complex telemetry environment of deep-space habitats, we use a modified version of the simulation model from [20] to generate degradation signals for engineered systems monitored by 20 sensors. Each system is assumed to experience one of two possible failure modes, FM 1 or FM 2, with only a subset of sensors being informative for degradation under each mode. This setup reflects the sensor redundancy and noise similar to onboard telemetry in a deep space habitat. We generate data for 200 systems per failure mode, using 160 systems for training and 40 for testing. To evaluate the robustness of our methodology, we simulate varying signal-to-noise ratios (SNRs) to mimic differences in sensor quality and environmental interference expected in long-duration missions. The resulting dataset includes multivariate sensor signals and the true time-to-failure (TTF), used for assessing prediction accuracy.
@@ 6.1 Simulation Model
The signals of the underlying degradation process for system $i$ exhibiting failure mode $k$ are generated according to Eq (22). The signal is parameterized by the random coefficient $\theta_{i}^{k}\sim N\left(\mu_{k},0.1^{2}\right)$, which is used to control the degradation rate. We let $\mu_{1}=1$, and $\mu_{2}=0.8$ be the degradation rates for FM 1 and FM 2, respectively. The TTF is computed as the first time the degradation signal $s_{i}^{k}(t)$ crosses a soft failure threshold $D_{k}$. We set the failure thresholds for FM 1 and FM 2 as $D_{1}=2$ and $D_{2}=1.5$, respectively.
@@
We consider a scenario where only a subset of sensors is correlated with the underlying degradation process associated with a given failure mode. These sensors are the informative sensors and their signal trends reflect the severity and progression of the degradation process. Note that sensors that do not exhibit any consistent trends leading up to a failure event are deemed noninformative. We define $\mathcal{J}_{k}$ as the index set of informative sensors for failure mode $k$, and $\mathcal{N}_{k}$ as the set of noninformative sensors such that $\mathcal{N}_{k}\cap\mathcal{J}_{k}=\{\}$ (null set). Out of the 20 sensors, we set $\mathcal{J}_{1}=\{5,12,16,19\}$ as the set of informative sensors for FM 1 and $\mathcal{J}_{2}=\{3,7,9,19\}$ as the set of informative sensors for FM 2. The $p^{th}$ sensor signal for the $i^{th}$ system exhibiting failure mode $k$ is generated using Eq (23).
@@
In this Eq, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the signal-to-noise for sensor $p$ and failure mode $k$. To account for the correlation between the signal trends of the informative sensors and the underlying degradation process, we generate $\theta_{i,p}^{k}$ from the following conditional distribution $\theta_{i,p}^{k}\mid\theta_{i}^{k}\sim N\left(\mu_{k}\left(1-\sqrt{1-\rho_{p}^{k}}\right),(0.1)^{2}\right)$ such that the correlation, $\rho_{p}^{k}$, is uniformly sampled from the interval $[0.80,0.99]$. The process for generating signals for the noninformative sensors is similar to that of the informative sensors, except the interval from which we sample $\rho_{p}^{k}$ is $[0.1,0.6]$. Finally, we randomly assign a sign to $\theta_{i,p}^{k}$ to allow for signals that increase or decrease in response to degradation.
@@ 6.2 Failure Mode Labeling and Sensor Selection
We analyze the ability of the MGR-ASGL model to accurately cluster the degradation signals into one of the two possible failure modes along with selecting the subset of informative sensors. We start by utilizing CA-FPCA as mentioned in 4.1 to extract features for each sensor separately. To determine the number of CA-FPC scores to retain, we use the 95% fraction of variance explained (FVE) criterion. For each cluster, we select the first $q_{p}^{k}$ CA-FPC-scores whose eigenvalues sum to at least 0.95. Since the number of CA FPC scores to retain can vary for each cluster, we retain the maximum of $q_{p}^{k}$ over both failure modes. For the remainder of this paper, we use 95% FVE criterion for determining the number of FPC scores, CA-FPC scores, and MFPC scores to retain. The CA-FPC scores serve as inputs to the MGR-ASGL model. Since this model has two tuning parameters, we utilize 5-fold cross-validation with $\lambda\in\{0.0050,0.0258,0.0466,...,0.3792,0.4000\}$ and $\alpha\in\{0.00,0.25,0.50,0.75,1.00\}$. We use minimum mean-squared error (minMSE) criterion, which selects the tuning parameters that minimize the mean-squared prediction error averaged over all folds. Table 2 displays the tuning parameter selection using both criteria over all three datasets. For each dataset and criterion, we fit the MGR-ASGL model using these tuning parameters. We then evaluate the performance of the MGR-ASGL model on the three datasets. While clustering is an unsupervised learning task, we know the ground truth regarding the failure modes. Therefore, we evaluate the ability of the clustering algorithm to match the ground truth. Furthermore, we have predetermined which sensors are most informative and we would like to determine whether or not the MGR-ASGL selects these sensors. The results of this performance evaluation are shown in Table 3.